Rick is 3 years older than half his brothers age. Rick is 13 how old is his brother?
Answer by stanbon(75887)

Rick is 3 years older than half his brothers age.
r = (1/2)b + 3
------
Rick is 13, how old is his brother?
------
13 = (1/2)b + 3
-----
(1/2)b = 10
-----
b = 20
--------
Cheers,
Stan H.
-----------